OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5).

OpenStudy (anonymous):

g(-5) = −2(-5) + 4 g(-5) = 10 + 4 g(-5) = 14

OpenStudy (anonymous):

?

OpenStudy (anonymous):

what's the problem?

OpenStudy (anonymous):

nothing someone commented saying porn..

OpenStudy (anonymous):

how strange..

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

yeah